OpenStudy (anonymous):

If the sides of a square are lengthened by 5cm, the area becomes 121cm^2. Find a length of a side of the original square.

OpenStudy (anonymous):

\[(x+5)^2= 121 ===> x=\sqrt(121) - 5 ====> x=7\]